Jefferson Lab HALL A Logbook for August 03 - All EXCEPT Automatic Entries

  • Main INDEX

    Sunday

    
    

    Saturday

  • 114287 08/30/03 22:39 Sulkosky Priapus Extracted
  • 114282 08/30/03 15:47 Sulkosky Target Tests III
  • 114281 08/30/03 15:38 Sulkosky EPR follow up
  • 114280 08/30/03 12:38 J. Singh fixed epr problem
  • 114279 08/30/03 00:59 Sulkosky Target Tests II
  • Friday

  • 114278 08/29/03 21:55 Sulkosky Target Tests and EPR Problems.
  • 114277 08/29/03 19:01 Sulkosky Target Tests I
  • 114276 08/29/03 14:06 Sulkosky Post Experiment Work
  • 114275 08/29/03 13:37 jpchen Septum training aborted deu to ESR 4k supply stopped
  • 114268 08/29/03 12:24 J. Singh last sGDH electron
  • 114267 08/29/03 12:11 bogdanw shift summary
  • 114214 08/29/03 08:41 J. Singh BWHP = IN
  • 114207 08/29/03 08:22 A. Tobias move to "Pol. 3He Cell" 0deg
  • 114202 08/29/03 08:08 J. Singh owl shift summary
  • 114198 08/29/03 08:05 J. Singh owl shift run summary
  • 114192 08/29/03 07:47 A. Tobias rotate target field & go to "Reference Cell"
  • 114189 08/29/03 07:25 A. Tobias rotate target field 0->270deg
  • 114184 08/29/03 07:02 J. Singh BHWP = OUT
  • 114177 08/29/03 06:50 A. Tobias moved to "Pol. 3He Cell" 0deg
  • 114172 08/29/03 06:31 A. Tobias "Reference Cell" w/ 140psig N2
  • 114167 08/29/03 06:23 A. Tobias/J. rotate field and cell change
  • 114160 08/29/03 05:55 J. Singh bleedthrough = 1.22%
  • 114157 08/29/03 05:45 A. Tobias rotate field from 0deg to 270deg
  • 114152 08/29/03 05:20 J. Singh BHWP = IN
  • 114145 08/29/03 04:56 A. Tobias NMR: 42.5% and EPR: none
  • 114137 08/29/03 04:10 J. Singh y_tg spectra for 1.977 GeV
  • 114134 08/29/03 04:08 A. Tobias moved to "Reference Cell"
  • 114123 08/29/03 03:43 J. Singh harp + otr looking good
  • 114122 08/29/03 03:41 A. Tobias "BeO-12C foils"
  • 114121 08/29/03 03:05 J. Singh harp scan looks wierd
  • 114120 08/29/03 02:27 J. Singh otr profile looks funny
  • 114116 08/29/03 00:57 A. Tobias EPR Wavetek 80 Output ON
  • 114115 08/29/03 00:26 A. Tobias target ladder is presently in "Empty Target" position & vent Ref. Cell
  • 114113 08/29/03 00:00 averett swing shift summary
  • Thursday

  • 114112 08/28/03 20:16 Yi Qiang NMR: 40.06% EPR: 39.02%
  • 114102 08/28/03 18:52 Sulkosky Reactz Spectra
  • 114097 08/28/03 18:36 Sulkosky ytg spectrum for Multi-Foil
  • 114096 08/28/03 18:34 Sulkosky Target Lifting Motor
  • 114067 08/28/03 16:25 Higinbotham Run Summary
  • 114066 08/28/03 16:25 Higinbotham Shift Summary
  • 114065 08/28/03 16:20 Higinbotham FSD with target changed to 1 foil
  • 114064 08/28/03 16:16 Higinbotham Checking Target
  • 114056 08/28/03 15:34 Ibrahim Target movements
  • 114032 08/28/03 13:16 Ibrahim Target movements
  • 114020 08/28/03 12:20 Ibrahim NMR = 38.7% and EPR = 37.4%
  • 114004 08/28/03 10:46 Ibrahim Field rotated 0 -> 270 deg
  • 113994 08/28/03 10:04 X. Jiang Run 4163, beam half wave plate is OUT.
  • 113982 08/28/03 09:23 Ibrahim Target changes
  • 113981 08/28/03 09:23 X. Jiang pol. he-3 run, at 0 deg.
  • 113965 08/28/03 08:18 Ibrahim Target changes
  • 113963 08/28/03 08:13 J. Singh owl shift summary
  • 113962 08/28/03 08:10 J. Singh owl shift run summary
  • 113961 08/28/03 08:09 J. Singh header files wrong for runs 4148 and 4149
  • 113959 08/28/03 07:59 J. Singh right runs 4150 and 4151 are junk
  • 113958 08/28/03 07:57 J. Singh left arm deadtime/prescale
  • 113955 08/28/03 07:46 A. Tobias rotate "Pol. 3He Cell" from 0deg to 270deg
  • 113953 08/28/03 07:42 A. Tobias NMR: 38.9% and EPR: 38.6%
  • 113944 08/28/03 06:40 J. Singh BHWP = IN
  • 113943 08/28/03 06:29 J. Singh bleedthrough = 0.91%
  • 113931 08/28/03 05:55 A. Tobias "Reference Cell" position discrepancy
  • 113924 08/28/03 05:45 A. Tobias move to "Pol. 3He Cell" at 0deg
  • 113919 08/28/03 05:25 A. Tobias fill "Reference Cell" w/ 142psig N2 gas
  • 113912 08/28/03 04:49 A. Tobias move to "Reference Cell" and rotate 3He target from 270deg to 0deg
  • 113906 08/28/03 04:30 A. Tobias back to "Pol. 3He Cell" at 270deg
  • 113905 08/28/03 04:24 J. Singh stripchart from low helium level in septum
  • 113904 08/28/03 04:21 A. Tobias Move to "Empty Target"
  • 113900 08/28/03 04:11 A. Tobias Rotated Target 0 to 270deg at 02:30
  • 113899 08/28/03 03:47 J. Singh MCC called: BCM trips
  • 113898 08/28/03 03:37 J. Singh BHWP = OUT
  • 113894 08/28/03 03:30 J. Singh bleedthrough = 1.87%
  • 113877 08/28/03 02:01 J. Singh BHWP = IN
  • 113862 08/28/03 01:13 A. Tobias move to "Pol. 3He Cell"
  • 113861 08/28/03 01:11 J. Singh bleedthrough measurement = 1.18%
  • 113858 08/28/03 01:04 A. Tobias move to "Reference Cell"
  • 113857 08/28/03 01:01 A. Tobias NMR: 40.3% EPR: 39.7%
  • 113849 08/28/03 00:01 E.Chudakov Shift summary
  • Wednesday

  • 113848 08/27/03 23:56 J,Yuan finish two runs of compton measurement :one is BHWP out,one is BHWP in
  • 113847 08/27/03 23:55 J,Yuan 23:55 rotate holding field from 270 to 0
  • 113793 08/27/03 20:57 jing,yuan rotate holding field from 0 to 270
  • 113772 08/27/03 18:20 Folts septum liquid level loss
  • 113770 08/27/03 17:54 Nanda Compton Result
  • 113769 08/27/03 17:45 Jing,Yuan NMR (41.7%)and EPR(41.0%)
  • 113743 08/27/03 16:06 jpchen Day shift summary
  • 113712 08/27/03 14:22 Ibrahim Target back to Pol. He
  • 113701 08/27/03 13:50 Sulkosky NMR: 43.2%, EPR: 41.4%
  • 113700 08/27/03 13:24 jpchen kin 11.10 completed. change to 11.11 and target back to 0 degree
  • 113687 08/27/03 12:26 jpchen BHWP out
  • 113669 08/27/03 11:46 Ibrahim Holding field rotated 0 -> 270 deg
  • 0 08/27/03 11:38 no_name no_key
  • 113666 08/27/03 11:36 jpchen target to 270 degree
  • 113665 08/27/03 11:35 jpchen beam half wave plate in
  • 113664 08/27/03 11:34 jpchen Bleedthrough=1.52%
  • 113652 08/27/03 11:14 jpchen left DAQ
  • 113645 08/27/03 10:42 jpchen production run resumed
  • 113642 08/27/03 10:30 gilman spot check
  • 113635 08/27/03 10:25 jpchen Compton tuning done, harp sacn, asked for raster on
  • 113634 08/27/03 09:48 jpchen Beam back, position good, Compton tuning
  • 113633 08/27/03 09:21 jpchen beam: MCC working on the separation
  • 113632 08/27/03 09:01 jpchen Beam tuning
  • 113630 08/27/03 08:12 J. Singh owl shift summary
  • 113629 08/27/03 07:49 J. Singh MCC called: sending tune beam
  • 113628 08/27/03 07:36 A. Tobias NMR: 44.9 % and EPR: ?
  • 113627 08/27/03 06:05 J. Singh MCC visit: no beam for ~ 2 hrs
  • 113626 08/27/03 05:03 J. Singh MCC visit: sending tune beam
  • 113625 08/27/03 02:57 J. Singh MCC visit: amplifier swap / redo separation
  • 113624 08/27/03 02:54 J. Singh MCC visit: RF separator cavity amplifier is dead
  • 113623 08/27/03 00:40 J. Singh MCC visit: RF separation problems
  • 113622 08/27/03 00:25 A. Tobias Empty Target
  • Tuesday

  • 113620 08/26/03 23:59 Jng,Yuan swing shift summary
  • 113619 08/26/03 23:48 Jing,Yuan MCC called beam will be back 1.50 hours later.they want to separate the beam with larger angle
  • 113618 08/26/03 23:15 jing,yuan MCC need more 15 minutes to tune
  • 113617 08/26/03 22:25 Jing,Yuan beam back to Hall A
  • 113616 08/26/03 21:29 Jing,Yuan MCC estimate beam back after 22:00pm
  • 113615 08/26/03 19:20 Jing,Yuan MCC called They are replacing all of the Arc 1 SCRs. Estimated time is 2 hrs.
  • 113614 08/26/03 18:09 Jing,Yuan rogram debuty Joe Grames called all the beam study will be concelled after today
  • 113613 08/26/03 18:06 Jing,Yuan ARC1 box supply problem from MCC
  • 113612 08/26/03 17:13 jpchen/jyuan pedestal runs
  • 113610 08/26/03 16:33 Jing ,Yuan MCC need more half an hour to send beam back to Hall A until now
  • 113608 08/26/03 16:10 bogdanw shift summary
  • 113607 08/26/03 13:10 Sulkosky EPR: 40.2%
  • 113606 08/26/03 13:08 A. Deur NMR, EPR and target to C12 foil
  • 113604 08/26/03 11:12 jpchen Left Q3 will be down for the rest of the running
  • 113603 08/26/03 11:09 A. Deur yt shoulder
  • 113602 08/26/03 11:09 jpchen target area inspection, colliamtors are fine
  • 113601 08/26/03 10:33 A. Deur Efficiency for kinematics 11.10
  • 113600 08/26/03 10:11 A. Deur run 2083 was Ne ref cell run
  • 113590 08/26/03 09:49 A. Deur Unusual empty target spectrum
  • 113586 08/26/03 09:37 Sulkosky Ytg Spectrum for Run # 4070
  • 113574 08/26/03 09:05 A. Deur Target polarization.
  • 113571 08/26/03 08:43 Sulkosky Kin. 11.10 Reference Cell Runs
  • 113560 08/26/03 08:03 R. Michaels owl shift summary
  • 113546 08/26/03 07:40 R. Michaels septum stuck 409, go for P = 3.104
  • 113544 08/26/03 07:23 J. Singh controlled access activities
  • 113543 08/26/03 07:11 R. Michaels septum was stuck
  • 113542 08/26/03 05:42 J. Singh target = empty target
  • 113541 08/26/03 05:37 R. Michaels access, set next kine 11.10
  • 113535 08/26/03 04:50 J. Singh target = ref cell => target = carbon multifoil
  • 113516 08/26/03 03:37 R. Michaels run 4071 - correction
  • 113509 08/26/03 03:33 J. Singh target = empty target
  • 113502 08/26/03 02:32 R. Michaels run 4069
  • 113496 08/26/03 01:44 R. Michaels current briefly to 10 uA
  • 113494 08/26/03 01:36 R. Michaels begin optics (kine 11.1)
  • 113489 08/26/03 01:15 R. Michaels R-arm magnets, Compton, etc
  • 113488 08/26/03 01:07 R. Michaels R-arm magnet controls problem
  • 113487 08/26/03 01:00 R. Michaels Compton retune done
  • 113481 08/26/03 00:07 Jing,Yuan swing shift summary
  • 113478 08/26/03 00:00 Jing,Yuan no_key
  • Monday

  • 113477 08/25/03 23:40 Folts/ Dezer septum JT
  • 113476 08/25/03 22:39 Sulkosky Septum He level Recovering
  • 113475 08/25/03 22:16 Eugene Moller results
  • 113474 08/25/03 22:04 Sulkosky Spectrometers Set
  • 113473 08/25/03 20:47 Sulkosky Right Dipole NMR not locking
  • 113472 08/25/03 20:12 Sulkosky Harp Scans after Compton Tuning
  • 113471 08/25/03 18:20 Jing,Yuan. WE finish OTR and Harps (x:199; y:126)
  • 113470 08/25/03 17:37 Jing,Yuan Call MCC beam back to check OTR and HARPS with 5uA.
  • 113469 08/25/03 17:29 Sulkosky Move target to Empty
  • 113468 08/25/03 17:23 jpchen left Q3 not working, should still take left arm data
  • 113467 08/25/03 17:09 Jing,Yuan power on Q1 and ramp up
  • 113466 08/25/03 16:55 jing,yuan Helium level of Q1(right arm) come back to 75%
  • 113464 08/25/03 16:02 jpchen, hers Day shift summary for 25 August 2003
  • 113463 08/25/03 16:01 jpchen Right dipole ramping back up, right Q1 problem!
  • 113462 08/25/03 15:25 P. Solvignon Laser alignment
  • 113461 08/25/03 14:59 jpchen target to pol. He3 position, field to 270 degree
  • 113460 08/25/03 14:50 jpchen controlled access
  • 113459 08/25/03 13:35 jpchen right dipole power supply problem
  • 113458 08/25/03 13:31 Sulkosky Septum Reached 418.25A
  • 113457 08/25/03 13:28 jpchen pass chaneg, beam tuning for the A line. 3 pm is estiamated time
  • 113456 08/25/03 12:21 J. Singh Right Dipole is wierd
  • 113455 08/25/03 12:07 J. Singh LQ1 & LQ3 liquid levels
  • 113454 08/25/03 12:04 A. Deur more on pions, pid and EMC effect
  • 113453 08/25/03 11:51 J. Singh runs 4048/24048 mislabeled
  • 113452 08/25/03 11:44 J. Singh septum current set incorrectly
  • 113451 08/25/03 11:26 O. Hansen data9 failure time; recovery unlikely
  • 113450 08/25/03 10:57 A. Deur 2-step process contamination for 4.4 GeV
  • 113449 08/25/03 10:37 Sulkosky Attempting Septum Training
  • 113448 08/25/03 10:26 jpchen Target collimators out, thick sieve in
  • 113447 08/25/03 10:10 J. Singh possibly issing runs were elastic
  • 113446 08/25/03 09:08 hersman, sul polarization 36.8%NMR 36.1%EPR
  • 113445 08/25/03 09:06 jpchen Data taking at kin 20 (4.4 GeV at 9 degree) completed. Septum training, and pass change (to 3 pass) started.
  • 113438 08/25/03 08:01 R. Snyder Owl Shift Target Summary
  • 113436 08/25/03 08:00 R. Michaels owl shift summary
  • 113426 08/25/03 06:56 R. Michaels ABU = Left. OR. Right DAQ
  • 113424 08/25/03 06:41 R. Michaels Automatic ABU fixed
  • 113420 08/25/03 06:24 J. Singh more on LQ3
  • 113416 08/25/03 06:14 J. Singh EPR = 37.24%/36.64%
  • 113415 08/25/03 06:14 R. Michaels automatic ABU wrong
  • 113412 08/25/03 05:42 R. Snyder NMR Measurement = 37.2%
  • 113391 08/25/03 03:24 R. Michaels He levels, called cryo., Q1 won't reset
  • 113383 08/25/03 02:38 R. Michaels Q1 L-arm trips off
  • 113353 08/25/03 00:30 R. Michaels missing runs due to data9 crash
  • 113349 08/25/03 00:04 Jing,Yuan swing shift summary
  • Sunday

  • 113343 08/24/03 23:58 A.Camsonne Target : reference cell filled with N2 140 psi
  • 113342 08/24/03 23:57 A. Camsonne Compton running
  • 113336 08/24/03 23:45 A. Camsonne Target moved to reference cell ( empty )
  • 113335 08/24/03 23:42 A. Camsonne Target polarization NMR : 37.32%
  • 113334 08/24/03 23:29 A. Camsonne Target moved to pick-up coils
  • 113333 08/24/03 23:23 jing,yuan Q3 of left arm quench
  • 113329 08/24/03 22:42 A. Camsonne Target : rotated holding field to 270 degres
  • 113328 08/24/03 22:35 jing,yuan rotate holding field from 0 to 270
  • 113318 08/24/03 21:59 Jing,Yuan we solve the problem of frenzen data window(reset the computer ,and restart)
  • 113317 08/24/03 21:42 Jing,Yuan Data window frenzed we cann't do anything
  • 113311 08/24/03 21:16 Sulkosky Pion/electron ratio for Kin. 12.17
  • 113300 08/24/03 20:43 Sulkosky Correction to Previous Entry
  • 113299 08/24/03 20:42 Sulkosky Ytg Spectrum for Kin. 12.16.5
  • 113298 08/24/03 20:35 Sulkosky Ytg Spectrum for Kin. 12.16.5
  • 113288 08/24/03 19:46 A. Camsonne Target moved to polarized He3 target
  • 113285 08/24/03 19:39 A. Camsonne Target reference cell filled with N2 140 psi
  • 113282 08/24/03 19:36 A. Camsonne Target moved to reference cell
  • 113279 08/24/03 19:28 Jing,Yuan spot++3.5/3.5
  • 113273 08/24/03 19:14 Jing,Yuan Septum change:I=234.8;P0=1.847
  • 113264 08/24/03 18:36 A. Camsonne Target moved to multifoil C target
  • 113263 08/24/03 18:35 A. Camsonne Target polarization NMR : 37.44 % EPR : 36.95 %
  • 113262 08/24/03 18:34 Jing,Yuan EPR(37.09%)/NMR(37.4%)
  • 113261 08/24/03 18:33 A. Camsonne Target moved to pick-up coils
  • 113260 08/24/03 18:22 Jing,Yuan septum change:P0=1.919Gev/c;I=244.4A
  • 113254 08/24/03 17:47 A. Camsonne Target : holding field to 270 degres
  • 113247 08/24/03 17:09 A. Camsonne Target polarization NMR : 37.51%
  • 113246 08/24/03 16:46 Jing,Yuan BHWP OUT
  • 113230 08/24/03 16:11 Lindgren shift summary
  • 113226 08/24/03 16:03 Lindgren kin 12.16
  • 113221 08/24/03 15:46 R. Michaels CODA revived again, clarified 'kcoda' instructions
  • 113214 08/24/03 15:16 Lindgren Kin 12.16
  • 113203 08/24/03 14:22 Lindgren kin 12.16
  • 113196 08/24/03 13:52 Lindgren Kin 12.16
  • 113193 08/24/03 13:45 R. Michaels CODA revived, /adaql2/data9 dead
  • 113181 08/24/03 12:45 Lindgren kin 12.16
  • 113178 08/24/03 11:50 Lindgren kin 12.15
  • 113174 08/24/03 11:15 hersman kinematics change to 12.16
  • 113163 08/24/03 10:26 hersman target polarization 36.58%
  • 113153 08/24/03 09:20 Lindgren Kin 12.5
  • 113150 08/24/03 09:16 hersman mouse problems
  • 113143 08/24/03 08:51 Lindgren kin 12.15
  • 113128 08/24/03 08:00 R. Michaels owl shift & run summary
  • 113126 08/24/03 07:59 R. Snyder Owl Shift Target Summary
  • 113125 08/24/03 07:58 R. Michaels dwell time of rastered spot
  • 113106 08/24/03 06:37 R. Michaels when R-arm DAQ hangs
  • 113085 08/24/03 05:20 R. Snyder NMR Measuremnet - 36.54%
  • 113073 08/24/03 04:47 R. Michaels PRL2 / 17 too
  • 113058 08/24/03 03:31 R. Michaels beam instability (was expected)
  • 113053 08/24/03 03:05 R. Michaels Possible detector problem - PRL2/15
  • 113042 08/24/03 02:15 R. Michaels spot size check
  • 113040 08/24/03 02:01 R. Snyder NMR Measurement = 36.51%
  • 113038 08/24/03 00:06 A. Deur Shift Summary
  • Saturday

  • 113037 08/23/03 23:38 J. Singh background correction spreadsheet
  • 113035 08/23/03 22:19 A. Deur Move target to empty position
  • 113034 08/23/03 22:13 Jing,Yuan move target to empty target
  • 113033 08/23/03 21:34 A. Deur Beam down. Maybe for one hour
  • 113030 08/23/03 21:04 Sulkosky Shower vs. Preshower
  • 113025 08/23/03 20:59 Sulkosky Pion Background increasing
  • 113022 08/23/03 20:53 A. Deur efficiency for kinematic 3.14 & rates
  • 113018 08/23/03 20:37 J. Singh EPR function generator
  • 113015 08/23/03 20:14 Sulkosky Target moved to Pol. 3He
  • 113007 08/23/03 20:04 Jing,Yuan fill the N2 into Ref. cell
  • 113005 08/23/03 19:58 J. Singh owl shift run summary 08/22
  • 113002 08/23/03 19:27 A. Deur quads cycling
  • 113001 08/23/03 19:26 J,Yuan NMR 33.3%
  • 112997 08/23/03 19:13 J. Singh run summaries
  • 112996 08/23/03 19:08 J. Singh run 3953 was ref cell = N2
  • 112988 08/23/03 18:50 JIng,Yuan rotate the holding field from 270 to 0
  • 112986 08/23/03 18:50 J. Singh corrected day shift run summary
  • 112973 08/23/03 18:11 A. Deur inconsistency in the replay
  • 112966 08/23/03 17:46 Jing,Yuan Move holding field from 0 to 270
  • 112963 08/23/03 17:37 Sulkosky Reactz for Mutli-12C foils
  • 112962 08/23/03 17:37 A. Deur Beam current
  • 112961 08/23/03 17:35 A. Deur Rate
  • 112958 08/23/03 17:24 A. Deur raster patern seems slightly skewed
  • 112953 08/23/03 17:16 A. Deur DAQ problem for left arm run 23959
  • 112948 08/23/03 16:55 A. Deur raster check
  • 112942 08/23/03 16:33 J. Singh day shift summary + bta
  • 112941 08/23/03 16:25 J. Singh day shift run summary
  • 112934 08/23/03 16:14 A. Deur Compton DAQ rebooted
  • 112910 08/23/03 14:18 Sulkosky EPR Measurement, P=32.3%
  • 112909 08/23/03 14:10 hersman beam pol'n 32.74
  • 112900 08/23/03 13:12 Hersman, Sul efficiency calculation spreadsheet
  • 112890 08/23/03 12:12 J. Singh missing beam time accounting
  • 112885 08/23/03 12:05 Hersman bleed through 1.82
  • 112882 08/23/03 11:45 B. Hersmen/S NMR: 31.85%
  • 112881 08/23/03 11:43 J. Singh BHWP = OUT
  • 112868 08/23/03 11:00 J. Singh bleedthrough = 1.82%
  • 112849 08/23/03 08:57 J. Singh MCC Called: beam back
  • 112848 08/23/03 08:32 R. Snyder Owl Shift Target Summary
  • 112847 08/23/03 08:22 J. Singh MCC called: RF zone dropped
  • 112846 08/23/03 08:12 J. Singh owl shift summary
  • 112841 08/23/03 08:05 J. Singh owl shift run summary
  • 112837 08/23/03 07:43 J. Singh BHWP = IN
  • 112826 08/23/03 06:38 J. Singh changing momentum to P0=2.704 GeV
  • 112819 08/23/03 05:25 J. Singh BHWP = OUT
  • 112807 08/23/03 04:40 R. Snyder NMR Measurement - 30.41%
  • 112782 08/23/03 01:50 J. Singh bleedthrough = 0.79%
  • 112777 08/23/03 01:44 J. Singh BHWP = IN
  • 112758 08/23/03 00:11 E.Chudakov Swing shift run summary
  • 112757 08/23/03 00:11 E.Chudakov Swing shift summary
  • Friday

  • 112752 08/22/03 23:52 Sulkosky Target Polarization Loss
  • 112739 08/22/03 23:08 yuan,jing NMR 28.0%
  • 112738 08/22/03 22:30 yuan,jing move target to Pol. 3He
  • 112730 08/22/03 21:59 Hersman move target position from 12C foils to reference cell
  • 112713 08/22/03 20:29 Sulkosky Target Strip Charts
  • 112705 08/22/03 19:53 Jing.Yuan NMR 25.2%
  • 112704 08/22/03 19:49 Nanda Compton Lock
  • 112703 08/22/03 19:48 Jing,Yuan Move target to Pol.3He
  • 112702 08/22/03 19:45 O. Hansen Day shift summary
  • 112689 08/22/03 19:14 Yuan,Jing move target to 12C multi-foils
  • 112683 08/22/03 18:31 Jing,Yuan EPR measurement .P=22.35
  • 112682 08/22/03 18:27 Jing,Yuan Pressure of N2 ref. cell
  • 112681 08/22/03 18:26 Sulkosky More on EPR
  • 112679 08/22/03 17:29 jpchen 12C foils replaced.
  • 112677 08/22/03 16:09 Yuan,Jing Move target from N2 reference cell to empty target
  • 112674 08/22/03 15:56 S. Bailey rotated holding field
  • 112659 08/22/03 14:40 kramerk Status of Pos. Fast Feedback
  • 112654 08/22/03 14:13 S. Bailey rotated holding field
  • 112640 08/22/03 12:58 J. Singh rough sketch of carbon foil state
  • 112614 08/22/03 10:52 Sulkosky ESPACE header file
  • 112611 08/22/03 10:21 S. Bailey NMR measurement
  • 112593 08/22/03 08:47 O. Hansen Septum stuck, HRSR to 3.1373 GeV
  • 112592 08/22/03 08:33 J. Singh owl shift summary
  • 112590 08/22/03 08:21 J. Singh controlled access activities
  • 112589 08/22/03 08:11 J. Singh Carbon Multifoil Damage
  • 112584 08/22/03 07:52 R. Snyder Owl Shift Target Summary
  • 112577 08/22/03 05:45 J. Singh 1.847 GeV/c ytg Spectrum
  • 112565 08/22/03 05:04 J. Singh 1.919 GeV/c ytg Spectrum
  • 112556 08/22/03 04:43 J. Singh beam position during 1.991 acceptance
  • 112553 08/22/03 04:30 J. Singh MCC called: BCM trips
  • 112542 08/22/03 04:07 J. Singh 1.991 GeV/c ytg Spectrum
  • 112539 08/22/03 04:01 J. Singh raster spot check + BCM trips
  • 112534 08/22/03 03:50 J. Singh OTR + harps scans
  • 112506 08/22/03 01:47 Eugene Moller results
  • 112504 08/22/03 00:00 E.Chudakov Swing shift summary
  • Thursday

  • 112503 08/21/03 23:22 E.Chudakov Swing shift runs summary
  • 112502 08/21/03 23:05 Nanda Compton Result
  • 112498 08/21/03 23:01 Sulkosky 2.704 GeV/c ytg Spectrum
  • 112495 08/21/03 22:55 Sulkosky Move target to Empty
  • 112488 08/21/03 22:31 Sulkosky 2.813 GeV/c ytg Spectrum
  • 112483 08/21/03 22:14 Sulkosky Move to Multi-foil 12C position
  • 112478 08/21/03 21:44 Sulkosky Target moved to Empty
  • 112477 08/21/03 21:43 Sulkosky 2.922 GeV/c Reference cell Runs
  • 112466 08/21/03 21:13 Sulkosky ytg spectrum for 2.992 GeV
  • 112459 08/21/03 21:02 Sulkosky Target moved to Empty
  • 112449 08/21/03 20:42 Sulkosky Target moved to Multi-Foil
  • 112440 08/21/03 20:14 E.Chudakov Harp scan
  • 112439 08/21/03 20:00 E.Chudakov Harp scan
  • 112438 08/21/03 16:06 Higinbotham Run Summary (day shift)
  • 112437 08/21/03 16:06 Higinbotham Shift Summary (day shift)
  • 112435 08/21/03 15:57 Rikki target shift summary
  • 112434 08/21/03 15:57 Higinbotham Septum Heating (part 2)
  • 112433 08/21/03 15:55 Higinbotham Septum Heating Test
  • 112432 08/21/03 15:47 Sulkosky Beam position strip chart
  • 112431 08/21/03 15:45 Sulkosky Beam Position Strip Chart
  • 112426 08/21/03 14:57 Higinbotham Current Calibration Data
  • 112421 08/21/03 14:51 Sulkosky Target Collimators removed
  • 112420 08/21/03 14:45 Sulkosky EPR Working!
  • 112408 08/21/03 12:14 J. Singh stripchart of temps
  • 112400 08/21/03 08:22 J. Singh lasers 5,6 were off!
  • 112397 08/21/03 08:11 R. Snyder Owl Shift Target Summary-Part 2
  • 112396 08/21/03 08:10 J. Singh owl shift summary part 2
  • 112392 08/21/03 08:03 J. Singh owl shift run summary part 2
  • 112383 08/21/03 07:18 R. Snyder NMR=39.06%
  • 112382 08/21/03 07:17 A. Puckett Q3, misbehaving
  • 112373 08/21/03 06:37 J. Singh BHWP = OUT
  • 112367 08/21/03 06:18 J. Singh beam spot diagnostics
  • 112366 08/21/03 06:11 J. Singh MCC called: beam back!
  • 112359 08/21/03 05:39 J. Singh owl shift summary, part 1
  • 112358 08/21/03 05:31 J. Singh owl shift run summary part 1
  • 112357 08/21/03 05:27 R. Snyder Owl Shift Target Summary-Part 1
  • 112356 08/21/03 05:20 J. Singh MCC called: update
  • 112355 08/21/03 04:39 J. Singh MCC called: problems, pt. 2
  • 112354 08/21/03 04:23 J. Singh MCC called: problems
  • 112352 08/21/03 04:04 J. Singh bleedthrough = 0.73%
  • 112333 08/21/03 02:28 J. Singh BHWP = IN
  • 112327 08/21/03 02:11 J. Singh bleedthrough = 0.25%
  • 112317 08/21/03 01:46 R. Snyder NMR =38.608%
  • 112314 08/21/03 01:16 A. Puckett Run 23830, left arm CODA reset
  • 112309 08/21/03 00:47 J. Singh BHWP = OUT
  • 112298 08/21/03 00:15 A. Deur shift summary part 2
  • 112294 08/21/03 00:08 Sulkosky Move to pol. 3He
  • 112293 08/21/03 00:05 Sulkosky Kin. 10.17 Reference cell runs
  • Wednesday

  • 112284 08/20/03 23:34 Sulkosky Target Changes
  • 112283 08/20/03 23:30 A. Deur done with kinematics 10.16
  • 112282 08/20/03 23:25 A. Deur correction to contamination entries
  • 112279 08/20/03 23:18 Sulkosky Rotate to 270 deg.
  • 112267 08/20/03 22:54 A. Deur insert bhwp.
  • 112265 08/20/03 22:53 A. Deur Contamination for 10:16
  • 112262 08/20/03 22:40 Sulkosky Move to pol. 3He
  • 112261 08/20/03 22:37 Sulkosky Kin. 10.16 Reference cell runs
  • 112252 08/20/03 22:07 Sulkosky Target Changes
  • 112251 08/20/03 22:06 A. Deur Done with kinematics 10.15.
  • 112249 08/20/03 21:50 Sulkosky Rotate to 270 deg.
  • 112243 08/20/03 21:38 A. Deur left/right run not synchronized
  • 112240 08/20/03 21:31 A. Deur insert beam 1/2 wave plate
  • 112239 08/20/03 21:29 A. Deur Bleedthrough measurement
  • 112235 08/20/03 21:15 A. Deur Contamination for 10:15
  • 112234 08/20/03 21:09 Punjabi spotsize
  • 112231 08/20/03 21:02 Sulkosky Move to pol. 3He cell
  • 112230 08/20/03 21:01 Sulkosky Reference cell runs
  • 112221 08/20/03 20:35 Sulkosky Target Changes, NMR = 39.4%
  • 112220 08/20/03 20:34 A. Deur shift summary, part 1
  • 112219 08/20/03 20:17 A. Deur Done with kinematics 10.14.
  • 112216 08/20/03 20:05 Sulkosky Rotate to 270 deg.
  • 112206 08/20/03 19:44 A. Deur left/right run not synchronized for run 3810/23809
  • 112203 08/20/03 19:39 A. Deur insert bhwp
  • 112198 08/20/03 19:38 Sulkosky ytg Spectrum from Kin. 10.11
  • 112197 08/20/03 19:37 A. Deur contamination for kinematics 10:14
  • 112192 08/20/03 19:26 Sulkosky Corrections
  • 112191 08/20/03 19:24 Sulkosky Move to Pol. 3He
  • 112188 08/20/03 19:19 Sulkosky Reference Cell Runs
  • 112182 08/20/03 18:59 Sulkosky Target Changes
  • 112177 08/20/03 18:34 A. Deur Kinematics 10:13 done
  • 112169 08/20/03 18:11 A. Deur left arm Q3 slightly instable
  • 112168 08/20/03 18:05 A. Deur beam half wave plat out
  • 112161 08/20/03 17:56 A. Deur raster check
  • 112157 08/20/03 17:44 Sulkosky Field Rotation to 270
  • 112154 08/20/03 17:38 A. Deur Hall C back: Our current dropped
  • 112151 08/20/03 17:32 A. Deur ask for Ibeam=8 uA
  • 112147 08/20/03 17:15 A. Deur left/right run not synchronized for run 3801/23800
  • 112141 08/20/03 17:06 A. Deur unexpected beam current increase
  • 112135 08/20/03 16:42 A. Deur Insert beam 1/2 wave plate
  • 112125 08/20/03 16:28 Sulkosky Beam positions not updating
  • 112118 08/20/03 16:04 R. Michaels Run summary part II
  • 112114 08/20/03 15:51 R. Michaels cron job to wake up NFS
  • 112113 08/20/03 15:49 jpchen NMR=40.17%
  • 112110 08/20/03 15:41 jpchen target to pol. He3, evacuated Ref cell.
  • 112107 08/20/03 15:34 jpchen Target to reference cell, empty then N2
  • 112099 08/20/03 15:06 R. Michaels raster spot size check
  • 112097 08/20/03 14:43 R. Michaels beam recovery is rocky
  • 112096 08/20/03 13:39 Sulkosky Controlled Access Details
  • 112095 08/20/03 13:00 jpchen correction: controlled access ended at 12:30
  • 112094 08/20/03 12:42 jpchen Controlled access for EPR ended (problem not fixed yet). Ready for beam.
  • 112093 08/20/03 11:31 A. Deur 2-step background for E=1.147 GeV
  • 112092 08/20/03 11:18 R. Michaels Q2 Left fixed / Magnets set / kine 10.13
  • 112091 08/20/03 11:12 R. Michaels day shift summary part I
  • 112090 08/20/03 11:12 Sulkosky EPR not possible!
  • 112089 08/20/03 11:06 Sulkosky NMR = 38.8%
  • 112072 08/20/03 09:11 jpchen Left Q2 polarity
  • 112063 08/20/03 08:37 R. Michaels switching Q2(L) polarity
  • 112052 08/20/03 08:01 J. Singh owl shift summary
  • 112051 08/20/03 07:58 J. Singh owl shift run summary
  • 112050 08/20/03 07:36 J. Singh MCC called: no beam ~ 15 min
  • 112047 08/20/03 07:12 Moffit Left Q2 NMR sign change?
  • 112046 08/20/03 07:05 Moffit MCC: No beam for 20 minutes
  • 112045 08/20/03 07:00 Moffit Drift in target coil currents
  • 112044 08/20/03 06:56 Moffit Target: Rotated holding field to 270°
  • 112043 08/20/03 06:50 J. Singh carbon target problem? pt 3
  • 112039 08/20/03 06:37 J. Singh carbon target problem? pt 2
  • 112034 08/20/03 06:26 J. Singh left arm carbon target problem?
  • 112025 08/20/03 05:50 Moffit Bleedthrough: 0.72%
  • 112018 08/20/03 05:29 Moffit BHWP IN
  • 111999 08/20/03 04:30 J. Singh bleedthrough = 1.03%
  • 111994 08/20/03 04:07 J. Singh MCC called: no beam ~ 10 min
  • 111991 08/20/03 03:59 Moffit Target moved to Pol. 3He
  • 111985 08/20/03 03:41 A. Puckett kinematic change
  • 111980 08/20/03 03:23 Moffit Target moved to Empty Reference Cell
  • 111979 08/20/03 03:22 Moffit Target: NMR = 38.5%
  • 111978 08/20/03 03:18 J. Singh bleedthrough = 2.99%->2.09%
  • 111966 08/20/03 01:43 A. Puckett right arm run-start failure
  • 111962 08/20/03 01:38 J. Singh BHWP = OUT
  • 111945 08/20/03 00:25 Moffit Charge Asymmetry Feedback Revived
  • 111938 08/20/03 00:01 Yuan,Jing swing shift summary
  • Tuesday

  • 111922 08/19/03 23:24 yuan,Jing left arm DAQ dead because TS1 don't report the status.WE reboot it again
  • 111921 08/19/03 23:12 X. Jiang target to 270 degree
  • 111913 08/19/03 22:49 Yuan,Jing HAPPEXE DIDN'T WORK
  • 111898 08/19/03 22:05 Yuan,Jing the Harp scan(H162;V 157) ,spot++(3.5/3.5),NMR measurement(38.3%)
  • 111891 08/19/03 21:36 X Jiang C12 single foil in
  • 111890 08/19/03 20:15 yuan,Jing Beam is ready to be back ;MCC just dod the final check
  • 111889 08/19/03 19:03 Yuan,Jing Until now,No beam .MCC guess Beam will be back after 20:00pm
  • 111888 08/19/03 17:35 Sulkosky NMR AFP Loss w/ Septum at 397.75A
  • 111887 08/19/03 17:20 Sulkosky Correction Coils and NMR!
  • 111886 08/19/03 16:57 Sulkosky EPR AFP Loss Tests
  • 111885 08/19/03 15:46 Sulkosky Target Tests with Septum
  • 111880 08/19/03 08:18 Rikki NMR & EPR
  • 111879 08/19/03 08:17 J. Singh owl shift summary
  • 111877 08/19/03 07:41 Sulkosky fixed right Q1 using red button
  • 111876 08/19/03 07:23 J. Singh spectrometers are set
  • 111875 08/19/03 06:58 J. Singh septum training time info
  • 111874 08/19/03 06:14 Moffit Right Q1 is not responding.
  • 111873 08/19/03 03:41 J. Singh status + owl shift summary , part 1
  • 111872 08/19/03 02:40 J. Singh controlled access activities
  • 111871 08/19/03 01:43 Moffit MCC: Probably no beam during OWL
  • 111870 08/19/03 01:18 J. Singh controlled access
  • 111869 08/19/03 00:56 Moffit Controlled Access
  • 111868 08/19/03 00:52 J. Singh magnet power supply faults
  • 111867 08/19/03 00:25 J. Singh lost He gas flow for 10 min in septum
  • 111866 08/19/03 00:11 Moffit Both Spectrometers have tripped
  • Monday

  • 111865 08/18/03 23:52 Yuan,Jing swing shift summary
  • 111864 08/18/03 23:44 Yuna,Jing septum situation
  • 111863 08/18/03 23:37 Yuan,jing 36 MCC called CHL have other problems ,so they don't know when beam will be back
  • 111862 08/18/03 19:32 Sulkosky Septum Temperature after Quench
  • 111861 08/18/03 19:29 Sulkosky Septum Quench at 427A
  • 111860 08/18/03 18:45 Holmstrom Step #2 of the BCM Linerarity Precedure
  • 111856 08/18/03 15:17 Holmstrom LHC Trip recovery Update
  • 111855 08/18/03 12:58 J. Singh quick elastic analysis @ 2.2 GeV (9 deg)
  • 111854 08/18/03 11:55 Sulkosky Target move to Empty position
  • 111853 08/18/03 11:49 R. Michaels day shift summary II
  • 111852 08/18/03 11:29 Sulkosky CHL Trip
  • 111851 08/18/03 11:15 J. Singh Elastic Runs @ 2.2 GeV / 9 deg
  • 111839 08/18/03 10:32 R. Michaels resume data taking (beam wave IN)
  • 111838 08/18/03 10:25 R. Michaels log entry delays related to NFS
  • 111837 08/18/03 10:25 jpchen NMR=41.7%, EPR=40.3%
  • 111833 08/18/03 10:08 R. Michaels partial day shift summary
  • 111824 08/18/03 09:43 R. Michaels halog test
  • 111799 08/18/03 08:27 Sulkosky Move to Pol. 3He
  • 111798 08/18/03 08:27 Sulkosky N2 Reference Cell Run
  • 111791 08/18/03 08:04 J. Singh owl shift summary
  • 111789 08/18/03 07:59 J. Singh owl shift run summary
  • 111784 08/18/03 07:30 Moffit Target move to reference cell
  • 111783 08/18/03 07:28 J. Singh right arm start run fails
  • 111773 08/18/03 06:47 J. Singh BHWP = OUT
  • 111772 08/18/03 06:45 J. Singh bleedthrough = 2.47%
  • 111765 08/18/03 06:36 Moffit Restart Xserver on adaqlr3
  • 111761 08/18/03 06:23 J. Singh MCC called: Fopt
  • 111755 08/18/03 06:14 Moffit Target: NMR = 42.3% + wierd
  • 111746 08/18/03 05:25 Puckett beam half-wave in/out
  • 111739 08/18/03 05:12 Puckett raster on/off?
  • 111734 08/18/03 05:05 J. Singh BHWP = IN
  • 111712 08/18/03 03:59 Moffit Target move to Pol. 3He
  • 111702 08/18/03 03:24 J. Singh bleedthrough = 2.82%
  • 111701 08/18/03 03:21 Moffit Target move to empty reference cell
  • 111698 08/18/03 03:19 J. Singh optics run summary, pt. 2
  • 111692 08/18/03 03:12 J. Singh beam coming back
  • 111689 08/18/03 02:31 Moffit MCC: RF Problems
  • 111681 08/18/03 01:38 Moffit Target move to carbon multi-foil
  • 111680 08/18/03 01:38 Moffit Target: NMR = 41.3%, EPR = 40.8%
  • 111652 08/18/03 00:00 Jing ,Yuan swing shift summary
  • Sunday

  • 111639 08/17/03 23:31 Yuan,Jing begin to new run for Pol3He cell,field 180,beam 1/2 plate OUT,run number(3706),Septum work stablely.
  • 111638 08/17/03 23:24 Yuan,Jing do harp scan (H120,V190),Spot++(3.5/3.5)
  • 111634 08/17/03 23:14 Yuan,Jing Beam back !!!!!!!!!!!!!!
  • 111630 08/17/03 22:34 Yuan,Jing Until now we still have no beam
  • 111624 08/17/03 19:40 Yuan,Jing MCC called other HV dropped , beam will be away half hou
  • 111612 08/17/03 19:09 Yuan,Jing beam back
  • 111610 08/17/03 19:05 Yuan,Jing two peaks in wire chamber for rigth arm ,one for left arm
  • 111609 08/17/03 18:27 Yuan,Jing beam away
  • 111603 08/17/03 17:53 Sulkosky NMR AFP loss w/ Septum @ 348.5A
  • 111593 08/17/03 17:15 comanl NMR measurement
  • 111589 08/17/03 16:46 Yuan,Jing Do bleedthrough 2.73%
  • 111584 08/17/03 16:31 Yuan,Jing beam position far away from target
  • 111571 08/17/03 16:01 A. Deur Run Summary
  • 111566 08/17/03 15:39 A. Deur bad beam
  • 111561 08/17/03 15:22 J. Singh Elastic Runs @ 1.1 GeV / 9 deg
  • 111554 08/17/03 14:54 R. Michaels prescale.dat needs a CR
  • 111553 08/17/03 14:52 R. Michaels halog numbering fixed
  • 111530 08/17/03 13:34 A. Deur Correction to end of run summary for runs 3680 to 3682
  • 111520 08/17/03 13:01 A. Deur Time and events for elastic runs
  • 111519 08/17/03 12:52 A. Deur Elastic runs
  • 111512 08/17/03 11:53 J. Singh correction to owl shift summary
  • 111507 08/17/03 11:28 A. Camsonne Target moved from C12 multifoils to polarized Helium 3
  • 111506 08/17/03 11:27 A. Deur optics runs done
  • 111501 08/17/03 11:15 A. Camsonne Target moved to C12 multifoils
  • 111500 08/17/03 11:15 J. Singh owl shift run summary
  • 111495 08/17/03 11:04 A. Camsonne Target : reference cell filled with N2 140 psig
  • 111492 08/17/03 10:59 A. Deur rate for empty ref cell
  • 111489 08/17/03 10:51 A. Camsonne Target moved to empty reference cell
  • 111486 08/17/03 10:44 A. Deur raster check
  • 111479 08/17/03 10:35 J. Singh left arm end of run summary
  • 111478 08/17/03 10:34 A. Camsonne Target moved to empty position
  • 111465 08/17/03 09:53 A. Camsonne Target polarization NMR : 42.6 % EPR : 40.3 %
  • 111464 08/17/03 09:51 A. Camsonne Counting house printer problem
  • 111463 08/17/03 09:48 A. Camsonne Target moved from pick-up coils to C12 Multifoil
  • 111462 08/17/03 09:45 A. Deur Collimators & frisker
  • 111461 08/17/03 09:39 J. Singh long lasers back on
  • 111460 08/17/03 09:29 A. Camsonne Target : laser off for collimator installation
  • 111459 08/17/03 09:22 A. Camsonne Target moved to pick-up coils
  • 111458 08/17/03 09:16 A. Deur HRS change + Access to put target collimator in.
  • 111448 08/17/03 08:37 A. Deur HRS at -3% but beam is down.
  • 111446 08/17/03 08:28 A. Deur raster check
  • 111445 08/17/03 08:28 A. Camsonne Target RTD screen
  • 111443 08/17/03 08:22 A. Camsonne Target moved from ref. cell to C12 Multifoil
  • 111439 08/17/03 08:12 A. Camsonne Target : Reference Cell filled with N2 138 Psi
  • 111432 08/17/03 08:01 J. Singh owl shift summary
  • 111431 08/17/03 07:56 Moffit Target move to empty reference cell
  • 111428 08/17/03 07:51 Moffit Beam Positions during Owl Shift
  • 111421 08/17/03 07:37 Moffit Target move to empty position
  • 111416 08/17/03 07:30 Moffit Left arm DAQ problems
  • 111413 08/17/03 07:26 J. Singh controlled access: sieve slit out
  • 111408 08/17/03 07:10 J. Singh bleedthrough = 2.05%
  • 111407 08/17/03 07:04 J. Singh bleedthrough = 2.15%
  • 111392 08/17/03 05:04 Moffit Target to carbon multi-foil
  • 111390 08/17/03 05:02 J. Singh left arm = 2.17 GeV
  • 111386 08/17/03 04:36 R. Feuerbach Adjusted HV settings on S2m
  • 111384 08/17/03 04:25 Moffit Target move to empty reference cell
  • 111381 08/17/03 04:16 J. Singh beam spot size
  • 111380 08/17/03 04:13 Moffit Target move to Empty position
  • 111376 08/17/03 04:04 J. Singh bleedthrough = 0.83%
  • 111373 08/17/03 03:54 J. Singh left arm = 2.06 GeV
  • 111371 08/17/03 03:45 J. Singh beam position
  • 111366 08/17/03 03:34 Moffit Target move to Carbon Multifoil
  • 111361 08/17/03 03:20 J. Singh HARP scans
  • 111358 08/17/03 03:02 j. cornejo raster spot size
  • 111345 08/17/03 01:39 Moffit Reboot pol3he3
  • 111343 08/17/03 00:23 Eugene Moller results
  • 111341 08/17/03 00:08 Jing,Yuan swing shift summary
  • Saturday

  • 111340 08/16/03 22:42 Sulkosky Move to Empty position
  • 111339 08/16/03 22:34 Sulkosky Clarification
  • 111338 08/16/03 22:29 Sulkosky Sieve pattern at 0%
  • 111337 08/16/03 21:59 Sulkosky Correction
  • 111336 08/16/03 21:57 Sulkosky 9 deg. Optics at 2.235 GeV
  • 111333 08/16/03 21:34 Sulkosky y_tg from run # 3643
  • 111321 08/16/03 21:07 Sulkosky 9 deg. Optics at 2.235 GeV
  • 111315 08/16/03 20:31 Yuan,J. Begin to do Optics
  • 111314 08/16/03 20:30 Sulkosky Move to Multi-foil position
  • 111313 08/16/03 20:13 Sulkosky Transverse Laser # 3
  • 111312 08/16/03 19:41 Sulkosky Controlled Access Activities
  • 111311 08/16/03 19:27 Yuan,J. harp scan
  • 111310 08/16/03 18:45 Yuan,J. have 2 passes set now
  • 111309 08/16/03 17:59 Sulkosky NMR: 39.11%, EPR: 38.48%
  • 111299 08/16/03 16:28 Sulkosky Target to Multi-Foil
  • 111291 08/16/03 16:05 Sulkosky Priapus Temperature Test # 6
  • 111289 08/16/03 16:03 bogdanw shift summary
  • 111274 08/16/03 15:16 S. Sirca DAQ on LHRS, xscalers for RHRS
  • 111237 08/16/03 13:00 comanl The target cannot be moved to polarized He cell -how to unstuck it
  • 111218 08/16/03 11:20 Sulkosky Priapus Temperature Tests # 4 and # 5
  • 111169 08/16/03 08:07 J. Singh owl shift summary
  • 111168 08/16/03 08:03 J. Singh owl shift run summary
  • 111165 08/16/03 07:51 J. Singh bleedthrough = 0.99%
  • 111164 08/16/03 07:50 Moffit Target move to Pol. 3He
  • 111157 08/16/03 07:19 Moffit Compton Lock off
  • 111153 08/16/03 07:06 Moffit Target: NMR= 40.2%, EPR=39.2%
  • 111147 08/16/03 06:17 Moffit Target: Rotate holding field to 270°
  • 111144 08/16/03 05:57 J. Singh bleedthrough = 0.75%
  • 111143 08/16/03 05:53 J. Singh bleedthrough = 0.43%
  • 111132 08/16/03 05:30 Moffit Target move to Pol. 3He
  • 111119 08/16/03 04:56 Moffit Left arm DAQ trouble...
  • 111117 08/16/03 04:37 Moffit Target Move to Empty Reference Cell
  • 111113 08/16/03 04:11 Moffit Target: Rotate holding field to 270°
  • 111104 08/16/03 03:25 Moffit Target Move to Pol. 3He
  • 111101 08/16/03 03:21 J. Singh checking beam spot
  • 111094 08/16/03 02:51 Moffit Move to Empty Reference Cell
  • 111091 08/16/03 02:11 Moffit Move to Empty Target
  • 111090 08/16/03 01:28 J. Singh MCC called: no beam
  • 111088 08/16/03 01:20 Moffit Beam Position Unacceptable
  • 111083 08/16/03 00:54 Moffit Target NMR: 39.7%
  • 111078 08/16/03 00:19 Moffit Target: Rotate holding field to 0°
  • 111077 08/16/03 00:11 J,Yuan swing shift summary
  • Friday

  • 111075 08/15/03 23:57 comanl We can't rotate the holding field
  • 111072 08/15/03 23:35 J,Yuan 23:35 We do the empty reference cell measurement again to correct We forget to vaculation refenrece cell before ,Just for Kin9.14 run
  • 111048 08/15/03 22:02 J,Yuan 20:11 we forget to vaculate N2 in reference cell,which maybe bring the large background for the Pol3He cell at Kinematics 9.13.We vacuum it at 21:50.
  • 111045 08/15/03 21:45 comanl Holding Field Rotation Display -wrong analog display
  • 111044 08/15/03 21:38 J,Yuan Gas cerenkov run number 3587,Kin9.13
  • 110981 08/15/03 16:59 potokar raster spot
  • 110974 08/15/03 16:29 jpchen 2-step process?
  • 110972 08/15/03 14:47 S. Sirca Field fluctuations in the LHRS dipole
  • 110971 08/15/03 13:10 K. McCormick Delay in returning from beam studies
  • 110970 08/15/03 11:39 Sulkosky new laser settings!
  • 110969 08/15/03 11:37 S. Bailey NMR & EPR
  • 110968 08/15/03 10:37 S. Bailey changed holding field
  • 110965 08/15/03 10:10 jpchen preshower/shower are fine
  • 110964 08/15/03 10:02 S. Bailey moved target
  • 110942 08/15/03 08:58 S. Bailey holding field rotated
  • 110922 08/15/03 08:03 peter Owl shift summary
  • 110918 08/15/03 07:54 J. Singh NMR = 40.4%, EPR = 39.4%
  • 110917 08/15/03 07:23 J. Singh moving to Pol 3He + vent ref cell
  • 110909 08/15/03 07:03 J. Singh ref cell = 100 psig N2
  • 110899 08/15/03 06:36 J. Singh "empty" ref cell y_tg spectra
  • 110892 08/15/03 06:29 Andrew Pucke Raster/Spot
  • 110883 08/15/03 06:00 J. Singh rotate 270 -> 0 & move to ref cell = empty
  • 110875 08/15/03 05:27 J. Singh rotating field 0 -> 270
  • 110870 08/15/03 05:10 J. Singh quick elastic analysis @ 1.1 GeV (9 deg)
  • 110856 08/15/03 03:56 J. Singh vent ref cell + moved to Pol 3He
  • 110853 08/15/03 03:53 J. Singh ref cell = 100 psig N2
  • 110851 08/15/03 03:51 J. Singh glitch during rotation 180->0
  • 110843 08/15/03 03:41 J. Singh new laser settings!
  • 110840 08/15/03 03:34 J. Singh 5,6,7 lasers on + ref cell = empty
  • 110839 08/15/03 03:31 J. Singh controlled access activities
  • 110835 08/15/03 03:23 J. Singh rotate to 0 deg
  • 110834 08/15/03 03:22 J. Singh nmr = 40.7%
  • 110814 08/15/03 01:27 peter preshower/shower
  • 110795 08/15/03 00:08 J,Yuan run number 3533/3534/3535/3536 detector data analyze finished
  • 110793 08/15/03 00:06 J,Yuan swing shift summary
  • 110790 08/15/03 00:02 comanl half wave plate in
  • 110787 08/15/03 00:00 J,Yuan ROC14 for left arm is on the state of disconnect,which let the left arm rate only 660hz
  • Thursday

  • 110786 08/14/03 23:58 J,Yuan We forget to change the new parameter for status of target 1/2 plate into DAQ.we had move into target half wave plate in before rotate holding filed from 0 to 180.So it apperared in end run number 3537,3538:target 1/2 plate OUT ,but should be really IN
  • 110778 08/14/03 23:17 comanl Rotate holding field from 0 to 180
  • 110774 08/14/03 23:08 J,Yuan there are 1 time each 30 min. alarm from the left arm septum dipole D-N,that lasted several seconds and disappeared
  • 110773 08/14/03 23:06 J,Yuan Detector data analyze
  • 110768 08/14/03 22:52 J,Yuan gas cerenkov #3,#4 low gain compared with #4,#6
  • 110735 08/14/03 21:10 Sulkosky Transverse Laser # 3
  • 110732 08/14/03 21:05 Sulkosky Target Collimators In
  • 110731 08/14/03 21:01 comanl NMR/EPR target measurements
  • 110707 08/14/03 18:19 Adaq Optics at 1.1 GeV
  • 110673 08/14/03 15:58 gilman day shift run summary
  • 110672 08/14/03 15:57 gilman day shift summary
  • 110671 08/14/03 15:42 Eugene Moller results
  • 110666 08/14/03 10:40 gilman target temp readback works on winbox
  • 110664 08/14/03 10:17 Wolfgang target polarization measurement
  • 110633 08/14/03 08:20 J. Singh left vs. right dipole fields
  • 110627 08/14/03 08:04 peter Owl shift summary
  • 110619 08/14/03 07:47 J. Singh left dipole alarm
  • 110605 08/14/03 06:52 peter Septum magnet strip chart
  • 110604 08/14/03 06:48 peter Right-arm VDC current draw
  • 110603 08/14/03 06:48 J. Singh controlled access activities
  • 110584 08/14/03 04:16 J. Singh changed alarm limit on cell RTD
  • 110561 08/14/03 02:40 peter Beam trips due to ion chambers
  • 110537 08/14/03 00:36 Higinbotham Shift Summary
  • Wednesday

  • 110535 08/13/03 23:32 Higinbotham Beam Back 11:25
  • 110534 08/13/03 19:02 Sulkosky NMR, EPR: 34.0%, 32.2%
  • 110533 08/13/03 18:06 Sulkosky Oven at 170 deg. C
  • 110532 08/13/03 17:40 Sulkosky RTD Temps on Windows
  • 110531 08/13/03 16:02 gilman day shift summary
  • 110529 08/13/03 15:29 Sulkosky Target Collimators removed, Transverse window replaced.
  • 110528 08/13/03 11:53 H.Yao 5 NMR Measurements
  • 110527 08/13/03 11:38 jpchen thin sieve inserted. ice checked.
  • 110526 08/13/03 11:03 H.Yao NMR & EPR
  • 110524 08/13/03 10:40 Wolfgang EPR result
  • 110523 08/13/03 09:53 gilman kin 8.26 done
  • 110520 08/13/03 09:36 gilman beam back
  • 110513 08/13/03 09:24 paschke spot size spec clarification
  • 110510 08/13/03 08:00 peter Owl shift summary
  • 110508 08/13/03 07:24 J,Yuan NMR 38.9%
  • 110507 08/13/03 06:58 peter Septum magnet chart
  • 110506 08/13/03 06:57 peter vacuum problem update
  • 110505 08/13/03 06:40 peter vacuum problem
  • 110496 08/13/03 05:36 peter Fast feedback on
  • 110476 08/13/03 04:15 peter kinematic change
  • 110475 08/13/03 04:09 peter Good Harp scan #2378
  • 110474 08/13/03 03:50 J,Yuan move target to single foil
  • 110469 08/13/03 01:29 peter No Fast Feedback
  • 110462 08/13/03 00:02 Higinbotham Shift Summary
  • Tuesday

  • 110460 08/12/03 19:38 J. Singh y target reconstruction check fo single foil
  • 110459 08/12/03 18:08 Sulkosky Longitudinal Lasers
  • 110455 08/12/03 14:25 Sulkosky Septum Quench
  • 110454 08/12/03 13:57 Vince increase septum ramp rate
  • 110453 08/12/03 12:26 Holmstrom Set Septum Current for quench
  • 110452 08/12/03 09:02 gilman septum quench
  • 110451 08/12/03 08:43 Wolfgang NMR and EPR measurements
  • 110448 08/12/03 08:01 B. Craver 8/12 Owl Shift Summary
  • 110442 08/12/03 07:30 Moffit Target: Rotate holding field to 270°
  • 110437 08/12/03 06:50 peter Correct septum magnet strip chart.
  • 110436 08/12/03 06:48 B. Craver Beam HWP moved IN
  • 110430 08/12/03 06:27 peter Septum magnet strip chart
  • 110427 08/12/03 06:19 Moffit Target to Pol. 3He
  • 110417 08/12/03 05:38 Moffit Target to reference cell
  • 110412 08/12/03 05:15 Moffit Target to Carbon Long Foil
  • 110406 08/12/03 04:47 Moffit Target: Rotate holding field to 270°
  • 110405 08/12/03 04:39 B. Craver Beam HWP moved OUT
  • 110397 08/12/03 04:23 B. Craver Beam HWP moved IN
  • 110396 08/12/03 04:17 Moffit BHWP In
  • 110381 08/12/03 03:28 Moffit Target to Pol 3He
  • 110372 08/12/03 02:59 peter Kinematic change to 8.23
  • 110371 08/12/03 02:55 Moffit Target to Reference Cell
  • 110366 08/12/03 02:33 Moffit Target to Carbon Long Foil
  • 110365 08/12/03 02:33 Moffit Target NMR: 39.5%
  • 110359 08/12/03 01:56 Moffit BHWP out
  • 110358 08/12/03 01:56 Moffit Target: Rotate holding field to 270°
  • 110350 08/12/03 01:25 Moffit BHWP IN
  • 110341 08/12/03 00:52 Moffit Move to Pol 3He
  • 110330 08/12/03 00:21 Moffit Target to Ref. Cell
  • 110328 08/12/03 00:15 Higinbotham Run Summary
  • Monday

  • 110320 08/11/03 23:59 Higinbotham Shift Summary
  • 110317 08/11/03 23:50 K. McCormick Kin 8.22 (Rotate target pol)
  • 110309 08/11/03 23:04 K. McCormick Rotated target field to 270 degrees
  • 110284 08/11/03 21:53 Higinbotham Bleedthru Check
  • 110271 08/11/03 21:17 Higinbotham Carbon Foil Check (kin 8.21)
  • 110265 08/11/03 21:05 P. Solvignon EPR Measurement
  • 110261 08/11/03 20:22 K. McCormick Rotated target polarization
  • 110230 08/11/03 18:57 Seonho Choi ESPACE update
  • 110223 08/11/03 18:35 paschke y_targ for Kin 8.20
  • 110201 08/11/03 17:24 K. McCormick Target changes
  • 110194 08/11/03 16:09 F. Butaru rotated holding field from o to 270 deg
  • 110193 08/11/03 16:08 Glashausser 8/11 Day shift summary
  • 110169 08/11/03 14:27 F. Butaru Moved target to Pol He3 position
  • 110168 08/11/03 14:26 F. Butaru Ref Cell was vented from 140 psig to 0 psig
  • 110164 08/11/03 14:07 F.Butaru Reference Cell filled in with N2 to 140 psig
  • 110151 08/11/03 13:21 Adaq magnet settings
  • 110150 08/11/03 13:19 F.Butaru Log summary
  • 110149 08/11/03 13:19 paschke y_targ for Kin 8.18
  • 110118 08/11/03 11:19 F.Butaru Rotaed the Holding Field from 0deg to 270 deg
  • 110117 08/11/03 11:18 glashausser Masking readback
  • 110104 08/11/03 10:16 glashausser Beam HWP
  • 110094 08/11/03 09:49 A. Deur run 3410 & 3411 are empty (no target) runs
  • 110087 08/11/03 09:22 F.Butaru Moved to Pol He3 position
  • 110069 08/11/03 08:12 Peter Owl shift summary
  • 110067 08/11/03 07:33 Moffit Septum ramped to 237.25
  • 110066 08/11/03 06:15 Moffit Septum ramped to 100A
  • 110065 08/11/03 06:03 Moffit Beam Permit
  • 110064 08/11/03 04:14 Moffit MCC estimate: No beam until Day Shift
  • 110063 08/11/03 02:59 Moffit NMR: 40.2%
  • 110062 08/11/03 01:44 A. Deur pions
  • 110060 08/11/03 00:07 B. Craver 8/10 Swing shift summary
  • Sunday

  • 110059 08/10/03 19:22 Sulkosky Reference Cell Position
  • 110058 08/10/03 18:54 A. Deur Ice in acceptance (?) & Collimator check
  • 110057 08/10/03 18:30 B. Craver N2 level for Q3 drops periodically
  • 110056 08/10/03 17:24 A. Deur cell windows are well into the acceptance
  • 110055 08/10/03 16:05 Glashausser Shift Summary 8/10/03
  • 110053 08/10/03 15:26 Glashausser ramp down septum
  • 110052 08/10/03 13:11 paschke CHL crash
  • 110039 08/10/03 12:22 P. Solvignon Cerenkov: no peak anymore at channel 1000
  • 110037 08/10/03 12:18 Wolfgang/Vin NMR and EPR
  • 110027 08/10/03 10:56 glashausser Beam HWP removed.
  • 110016 08/10/03 10:29 A. Deur Cerenkov additional peak & pions
  • 110013 08/10/03 10:27 glashausser CODA goal calculation
  • 110012 08/10/03 10:08 glashausser Beam HWP
  • 109996 08/10/03 09:06 paschke y-targ for Kin 8.15
  • 0 08/10/03 09:00 no_name no_key
  • 109984 08/10/03 08:34 paschke y-targ for Kin 8.15
  • 109979 08/10/03 08:05 Moffit Owl Shift Summary
  • 109978 08/10/03 08:03 Moffit HRS/Septum Settings for 8.15
  • 109976 08/10/03 07:49 H.Yao Rotate Field from 270 to 0 and Move target to Reference Cell
  • 109967 08/10/03 06:59 H.Yao Rotate Holding Field from 0 to 270
  • 109966 08/10/03 06:59 Moffit BHWP Out
  • 109958 08/10/03 06:21 Moffit Bleedthrough: 1.6%
  • 109953 08/10/03 05:58 Moffit BHWP IN
  • 109934 08/10/03 05:02 Moffit End of run failed: 23386
  • 109928 08/10/03 04:51 Moffit Y-target for 8.14. Carbon
  • 109925 08/10/03 04:34 H.Yao Do NMR Measurement
  • 109924 08/10/03 04:31 Moffit RHRS, Septum settings for 8.14
  • 109923 08/10/03 04:16 H.Yao Rotate Field to 0 and move target to Single Foil 12C
  • 109915 08/10/03 03:34 Moffit Septum Stripchart
  • 109912 08/10/03 03:16 H.Yao Rotate Holding Field from 0 to 270
  • 109909 08/10/03 03:06 Moffit Re: Gas Cerenkov
  • 109906 08/10/03 02:53 Moffit Spike in GASADCR
  • 109905 08/10/03 02:45 R. Michaels test halog (log # file was missing)
  • Saturday

  • 109856 08/09/03 23:46 Sulkosky y_tg Spectrum for Kin 8.13
  • 109843 08/09/03 22:57 Sulkosky Rotate to 0 deg.
  • 109842 08/09/03 22:49 Sulkosky Rotate to 270 deg.
  • 109832 08/09/03 21:59 B. Craver Beam HWP moved OUT
  • 109820 08/09/03 21:19 B. Craver Beam HWP moved IN
  • 109819 08/09/03 21:08 Sulkosky Reference cell vented
  • 109814 08/09/03 21:01 Sulkosky Cut adjustments in Replay
  • 109811 08/09/03 20:47 Sulkosky Kin 8.12, y_tg spectrum
  • 109809 08/09/03 20:36 Sulkosky y_tg at 324.5A on Septum
  • 109790 08/09/03 19:30 Sulkosky NMR;EPR = 35.8%;34.0%
  • 109789 08/09/03 19:10 Sulkosky Rotate to 0 degrees
  • 109788 08/09/03 19:02 Sulkosky Rotate to 270 degrees
  • 109773 08/09/03 17:57 B. Craver pion rejector detector replay
  • 109761 08/09/03 17:06 B. Craver Beam HWP moved OUT
  • 109750 08/09/03 16:29 Glashausser 8/9/03 day run summary
  • 109722 08/09/03 15:11 paschke RHWP scan aborted
  • 109721 08/09/03 15:10 J,Yuan Move target to empty reference cell
  • 109714 08/09/03 14:39 glashausser BPM change with Hall C on/off
  • 109713 08/09/03 14:38 Sulkosky Y_tg at 348.5A on Septum
  • 109709 08/09/03 14:25 glashausser Septum calibration constant changed
  • 109688 08/09/03 13:10 J,Yuan detector/physics data analyze :run number 3333,3334,3335,3336.Det. work well
  • 109686 08/09/03 13:09 J,Yuan Move target to single foil
  • 109683 08/09/03 12:37 Sulkosky Pick-up Coils Position II
  • 109680 08/09/03 12:31 paschke bleedthrough measurement
  • 109677 08/09/03 12:23 Sulkosky Pick-up Coils Position
  • 109667 08/09/03 11:28 J,Yuan NMR measurement 36.06%
  • 109666 08/09/03 11:23 paschke Kin 8.10 single foil, y_target
  • 109654 08/09/03 10:11 J,Yuan Move target to Pol. 3He
  • 109641 08/09/03 08:57 J,yuan Move target to empty target because the beam position drifts
  • 109624 08/09/03 08:00 B. Moffit Owl Shift Summary
  • 109591 08/09/03 06:46 F.Butaru Moved target to Polarized He3 position
  • 109590 08/09/03 06:45 F. Butaru Moved to Be+12C Long Foil position
  • 109589 08/09/03 06:27 Moffit MCC: Beam away to check orbits
  • 109583 08/09/03 06:16 F.Butaru NMR Polarization Measurement
  • 109581 08/09/03 06:09 Moffit RHRS/Septum Settings for 8.10
  • 109580 08/09/03 06:08 F. Butaru Pick-Up Coils Position wasn't displayed corectley
  • 109566 08/09/03 05:02 Moffit RightArm DAQ problem
  • 109559 08/09/03 04:50 Moffit BHWP In
  • 109542 08/09/03 04:25 Moffit Spec_R.P in physics replay
  • 109533 08/09/03 04:07 F. Butaru Rotated Field direction to 270 deg
  • 109528 08/09/03 03:40 Moffit BHWP Out
  • 109505 08/09/03 03:00 Moffit BHWP In
  • 109465 08/09/03 01:51 F. Butaru Pol 3He on position
  • 109464 08/09/03 01:50 F. Butaru N2 Ref Cell vented
  • 109461 08/09/03 01:40 F. Butaru Reference Cell filled with N2
  • 109456 08/09/03 01:29 F. Butaru Moved to Empty Reference Cell position
  • 109452 08/09/03 01:25 Moffit Pockels Cell Voltage change
  • 109451 08/09/03 01:25 F. Butaru Moved to Empty Target position
  • 109445 08/09/03 01:07 Sulkosky NMR = 38.4%, EPR = 35.3%
  • 109444 08/09/03 01:02 jpchen Thick sieve removed, target collimators installed, ice removed.
  • 109443 08/09/03 00:17 Moffit Controlled Access
  • Friday

  • 109441 08/08/03 23:51 B. Craver N2 ref. cell
  • 109440 08/08/03 23:50 jpchen Swing shift syummary
  • 109429 08/08/03 23:22 Margaziotis Raster/Empty tgt
  • 109410 08/08/03 22:49 Sulkosky Septum Tune at 9 degrees
  • 109406 08/08/03 22:46 jpchen Beam size fine now
  • 109405 08/08/03 22:41 jpchen optics tuning
  • 109403 08/08/03 22:39 Sulkosky More Septum Tuning
  • 109391 08/08/03 22:04 Sulkosky Septum Tuning at 9 degrees
  • 109364 08/08/03 20:44 jpchen Compton tuning: electron detector not moving
  • 109361 08/08/03 19:57 jpchen spot size
  • 109358 08/08/03 19:53 Eugene Moller results
  • 109357 08/08/03 18:37 paschke RHWP in source / new PC voltages
  • 109356 08/08/03 18:21 Sulkosky Run Table for 3.775 GeV
  • 109352 08/08/03 17:33 jpchen Bema spot too small. Moller will start now
  • 109351 08/08/03 17:25 jpchen beam back. Beam tuning.FF not working
  • 109347 08/08/03 15:43 J,Yuan Detector data analyze(cosmic ray) run number 3264,3265:right arm work well
  • 109346 08/08/03 15:27 Lindgren Compton polarimeter stll not working
  • 109345 08/08/03 14:36 Lindgren Hall A access
  • 109344 08/08/03 14:35 Lindgren Harp scan
  • 109343 08/08/03 14:19 paschke parity DAQ: lockfile trick
  • 109342 08/08/03 14:16 glashausser spot size
  • 109340 08/08/03 14:06 Lindgren Beam spot size check
  • 109327 08/08/03 13:06 Lindgren Beam back
  • 109326 08/08/03 10:49 Lindgren Hall A access
  • 109325 08/08/03 10:39 Lindgren shift comments
  • 109324 08/08/03 09:17 Lindgren shift comments
  • 109323 08/08/03 08:43 Lindgren shift comments
  • 109322 08/08/03 08:03 Holmstrom Owl Shift Summary
  • 109320 08/08/03 07:43 A. Shabetai RTDs temperatures
  • 109319 08/08/03 07:39 Sulkosky EPR: 30.0%
  • 109316 08/08/03 06:24 A. Shabetai NMR = 30.24 %
  • 109315 08/08/03 05:56 Holmstrom Trim Rack out of Commision
  • 109314 08/08/03 05:44 Holmstrom Ion Chamber trouble
  • 109311 08/08/03 05:09 Holmstrom HV trip again
  • 109310 08/08/03 04:19 Holmstrom HV trip
  • 109306 08/08/03 03:34 Holmstrom Beam Update
  • 109305 08/08/03 03:31 Holmstrom HV update
  • 109304 08/08/03 03:15 Holmstrom HV update
  • 109303 08/08/03 03:12 Holmstrom Left Arm Q3 current alarm
  • 109302 08/08/03 02:02 A. Shabetai Laser Spectrum is low
  • Thursday

  • 109300 08/07/03 23:23 G. Chang Septum magnet
  • 109299 08/07/03 21:17 G. Chang Beam to the hall
  • 109298 08/07/03 18:45 paschke warm He consumption
  • 109297 08/07/03 18:40 jpchen right magnet all set
  • 109296 08/07/03 18:13 Sulkosky Longitudinal Laser #5 On
  • 109295 08/07/03 18:09 Sulkosky EPR: 17.05%
  • 109294 08/07/03 17:46 Sulkosky Reference cell Encoder Position
  • 109293 08/07/03 16:19 G. Chang Lasers on
  • 109292 08/07/03 15:43 K. McCormick Lasers off
  • 109287 08/07/03 15:28 jpchen Left magnets are set
  • 109286 08/07/03 15:24 jpchen pedestal runs
  • 109284 08/07/03 15:06 jpchen MD work completed in Hall A, sweep in progress
  • 109283 08/07/03 14:45 jpchen added another 12C at +20 cm
  • 109282 08/07/03 14:36 Sulkosky Septum Training
  • 109281 08/07/03 13:15 R. Feuerbach Left HRS S2m HV adjusted
  • 109275 08/07/03 11:23 jpchen Added a short 12C at -10 cm for multi-12C target
  • 109258 08/07/03 04:46 Holmstrom Ramp the Septum
  • Wednesday

  • 109255 08/06/03 20:56 P. Solvignon Laser alignment
  • 109245 08/06/03 17:22 jpchen Maintenance activities (II)
  • 109244 08/06/03 16:47 Seonho Choi 9 degree database
  • 109205 08/06/03 10:32 jpchen Maintenance work (I)
  • Tuesday

  • 109203 08/05/03 18:44 Jack CO2 gas system
  • 109202 08/05/03 18:14 slifer Compensating coils
  • 109198 08/05/03 10:26 paschke halog web server down
  • 109197 08/05/03 09:30 jpchen lasers locked out
  • 109196 08/05/03 09:04 Sulkosky Turn Oven Off
  • 109195 08/05/03 08:57 Sulkosky Septum Quench at 412.5A
  • 109194 08/05/03 08:52 Sulkosky Septum Training
  • 109192 08/05/03 08:18 Sulkosky Lasers turned off
  • 109191 08/05/03 08:00 K. Keister cooling jets off
  • 109190 08/05/03 07:43 R. Feuerbach Problems with S2m histograms in ESPACE
  • 109189 08/05/03 07:39 K. Keister owl shift target summary
  • 109188 08/05/03 06:09 Sulkosky Run Summary
  • 109187 08/05/03 06:07 Sulkosky Shift Summary I
  • 109174 08/05/03 05:21 Sulkosky Bleedthrough
  • 109169 08/05/03 05:16 Sulkosky Bleedthrough Measurements
  • 109120 08/05/03 01:07 Sulkosky Bleedthrough: 2.59%
  • 109106 08/05/03 00:22 Reitz swing shift summary II
  • Monday

  • 109094 08/04/03 23:29 Reitz septum temps
  • 109064 08/04/03 21:52 Reitz DataCube 1H04A
  • 109047 08/04/03 20:30 Reitz beam half wave plate in
  • 109019 08/04/03 17:41 Reitz medm problems
  • 109018 08/04/03 17:26 reitz beam half wave plate out
  • 108996 08/04/03 16:04 jpchen bleedthrough 2.69%
  • 108995 08/04/03 16:02 jpchen Day Shift Summary
  • 108989 08/04/03 15:43 J,Yuan Turn off the transverse lasers and rotate holding field from 270 to 0
  • 108988 08/04/03 15:28 jpchen Changing to kin 7.15
  • 108978 08/04/03 14:55 jpchen beam back
  • 108977 08/04/03 13:29 jpchen beam away
  • 108976 08/04/03 13:13 J,Yuan Turn on transverse lasers
  • 108975 08/04/03 13:09 J,Yuan rotate holding field from 0 to 270;
  • 108967 08/04/03 12:43 jpchen BHWP in
  • 108946 08/04/03 11:51 J,Yuan Move target to Pol. 3He cell
  • 108941 08/04/03 11:38 jpchen 140 psig N2
  • 108935 08/04/03 11:14 jpchen changed to kin 7.14
  • 108934 08/04/03 11:12 jpchen NMR=41.69%, EPR=40.50%
  • 108933 08/04/03 11:00 paschke Moved to Kinematics 7.14
  • 108920 08/04/03 10:21 adaq BHWP out
  • 108911 08/04/03 09:52 jpchen pol, He3 0 degree for kin 7.13
  • 108903 08/04/03 09:17 R. Michaels end-run scripts
  • 108898 08/04/03 09:12 jpchen kin 7.13, pol 270 completed. MOved to Empty ref
  • 108887 08/04/03 08:39 Sulkosky Septum Cool Down
  • 108886 08/04/03 08:09 Sulkosky Run Summary
  • 108881 08/04/03 08:02 K. Keister owl shift target summary
  • 108880 08/04/03 08:00 Sulkosky Shift Summary
  • 108863 08/04/03 07:15 paschke Reviving Compton Polarimeter
  • 108860 08/04/03 06:54 paschke Charge feedback revived
  • 108840 08/04/03 05:46 K. Keister spot++ raster check
  • 108833 08/04/03 03:37 Sulkosky Beam Orbits bad
  • 108832 08/04/03 03:01 J. Singh curiosity in begin / end of run summary
  • 108831 08/04/03 02:52 Sulkosky Run Table for 2.845 GeV
  • 108830 08/04/03 02:28 J. Singh Elastic Runs @ 2.8GeV/6 deg
  • 108829 08/04/03 02:07 Sulkosky Beam Tuning
  • 108828 08/04/03 00:58 Sulkosky Beam estimate: 3 am
  • 108825 08/04/03 00:03 P. Solvignon Shift summary
  • Sunday

  • 108821 08/03/03 23:13 P. Solvignon Beam delivery time estimate: 2:00am
  • 108820 08/03/03 23:11 micha niskin moved target to empty target position
  • 108819 08/03/03 22:46 P. Solvignon All cryo levels back up, magent currents ramping up
  • 108818 08/03/03 21:19 J. Singh septum stripchart last 11 hrs
  • 108817 08/03/03 21:14 J. Singh HRS update
  • 108816 08/03/03 21:07 J. Singh septum update
  • 108815 08/03/03 17:44 micha niskin reboot of labview machine
  • 108814 08/03/03 17:39 P. Solvignon Septum is cooling down
  • 108813 08/03/03 17:18 micha niskin moved target to pickup coil position
  • 108811 08/03/03 15:56 J,Yuan Day shift summar II
  • 108810 08/03/03 15:37 J,Yuan CHL Reports that ESR will be ready to take loads to Halls in 1hr
  • 108809 08/03/03 14:54 J,Yuan 14:45 Septum situation
  • 108808 08/03/03 14:52 J,Yuan magnet cryogen levers
  • 108807 08/03/03 14:13 J,Yuan Septum's temp. rise so fastly and the Cool Helium still don't come back
  • 108806 08/03/03 13:36 J,Yuan Septum continue to be warmer
  • 108805 08/03/03 13:04 J,Yuan Day shift summary I
  • 108804 08/03/03 12:54 J. Singh ESR tripped, again
  • 108803 08/03/03 11:14 paschke ESR crash
  • 108802 08/03/03 11:11 J,Yuan HRS CRASHED AGAIN
  • 108801 08/03/03 11:09 J,Yuan HRS crashed again
  • 108799 08/03/03 07:58 Sulkosky Shift Summary III
  • 108798 08/03/03 04:36 Sulkosky EPR-NMR Calibration
  • 108795 08/03/03 04:04 Sulkosky Hall A Tools
  • 108794 08/03/03 03:39 Sulkosky Shift Summary II
  • 108793 08/03/03 01:24 Sulkosky Shift Summary I
  • 108791 08/03/03 00:11 B. Craver shift summary
  • Saturday

  • 108790 08/02/03 17:06 J. Singh Hall C lost HMS at same time
  • 108789 08/02/03 16:48 Lhuillier Zero Compton asym
  • 108788 08/02/03 16:43 J. Singh Left Arm Q3 N2 level recovery
  • 108787 08/02/03 16:35 J. Singh Right arm HRS power supply faults
  • 108786 08/02/03 16:35 J. Singh Left Arm HRS power supply faults
  • 108785 08/02/03 16:22 J. Singh Left Q3 Liquid Nitrogen Level
  • 108783 08/02/03 16:05 J. Singh HRS Crash
  • 108782 08/02/03 16:00 J. Singh Septum LHe level dipped
  • 108781 08/02/03 15:59 H. Gao Shift run summary
  • 108780 08/02/03 15:45 J,Yuan right arm detector status
  • 108779 08/02/03 15:43 J,Yuan Move target to empty refernce cell;
  • 108778 08/02/03 15:16 J,Yuan begin to do the longitude measurement
  • 108776 08/02/03 14:59 R. Michaels Parity "GDH" works now.
  • 108775 08/02/03 14:57 R. Michaels L-arm DAQ ok now, network pbm fixed
  • 108769 08/02/03 14:11 H. Gao CHL problem
  • 108767 08/02/03 13:11 H. Gao lost beam
  • 108739 08/02/03 10:03 H. Gao Septum setting change
  • 108731 08/02/03 09:11 H. Gao run 3174 & 3175
  • 108726 08/02/03 08:57 H. Gao beam half-wave plate out
  • 108714 08/02/03 08:05 S. Woo laser : transvers, holding field : 270, THWP : OUT
  • 108713 08/02/03 08:03 J. Singh owl shift summary
  • 108711 08/02/03 08:00 J. Singh owl shift run summary
  • 108705 08/02/03 07:23 J. Singh BHWP = IN
  • 108704 08/02/03 07:22 J. Singh bleedthrough = 2.1%
  • 108699 08/02/03 07:08 R. Stringer Run start header incomplete
  • 108694 08/02/03 06:41 S. Woo Holding field : 180 , THWP : IN
  • 108692 08/02/03 06:27 J. Singh BHWP = OUT
  • 108688 08/02/03 06:01 J. Singh controlled access + DAQ problem, part 2
  • 108685 08/02/03 05:49 R. Stringer Left arm angle readback
  • 108673 08/02/03 05:00 R. Michaels Parity DAQ -- use GDH_Minimal
  • 108669 08/02/03 04:45 S. Woo NMR : 40.45%
  • 108668 08/02/03 04:37 R. Michaels (DAQ) adaqs2 192 subnet problem
  • 108666 08/02/03 03:46 R. Michaels DAQ problems - adaqs2
  • 108665 08/02/03 03:19 J. Singh DAQ problem, part 1
  • 108663 08/02/03 02:25 J. Singh quick elastic analysis @ 2.8GeV
  • 108662 08/02/03 01:40 S. Woo Pol 3he cell
  • 108658 08/02/03 01:19 S. Woo filled N2 in Ref. Cell
  • 108653 08/02/03 00:52 S. Woo EPR : 38.09% , empty ref. cell
  • 108652 08/02/03 00:33 S. Woo NMR : 38.26%
  • 108647 08/02/03 00:03 E.Chudakov Swing shift summary
  • Friday

  • 108605 08/01/03 22:17 micha niskin rotated holding field to 270 deg
  • 108579 08/01/03 20:36 micha niskin rotated holding field to 0 deg
  • 108550 08/01/03 19:00 micha niskin moved target to polarized 3He position
  • 108538 08/01/03 17:06 micha niskin filled reference cell with N2, 140 psig
  • 108527 08/01/03 16:17 micha niskin moved target to reference cell
  • 108525 08/01/03 16:02 gilman bleedthrough measurement
  • 108524 08/01/03 16:01 gilman day shift summary, August 1
  • 108523 08/01/03 15:58 gilman day shift run summary
  • 108498 08/01/03 14:34 J. Singh Neutron Spin Crisis
  • 108497 08/01/03 14:30 S. Bailey 1/2 wave plate correction
  • 108491 08/01/03 14:20 micha niskin NMR = 37.6
  • 108460 08/01/03 11:45 J,Yuan move the target to Pol 3He cell
  • 108459 08/01/03 11:24 paschke New PC-HV (PITA) setpoints
  • 108449 08/01/03 11:05 J,Yuan 11:00 Filled in N2 reference cell (140Psig)
  • 108437 08/01/03 10:21 J,Yuan move target to empty ref. cell
  • 108433 08/01/03 10:02 paschke PITA scan, Charge Asymmetry controlled
  • 108417 08/01/03 08:53 Reitz beamline HV
  • 108404 08/01/03 08:04 J. Singh owl shift summary
  • 108402 08/01/03 08:01 J. Singh owl shift run summary
  • 108401 08/01/03 07:56 J. Singh compton + eP HV: something wrong?
  • 108396 08/01/03 07:45 J. Singh bleedthrough = 1.07%
  • 108395 08/01/03 07:44 J. Singh BHWP = OUT
  • 108383 08/01/03 07:07 J. Singh bleedthrough = 1.36%
  • 108368 08/01/03 06:11 J. Singh optics @ 2.844GeV run summary
  • 108356 08/01/03 05:24 J. Singh EPR = 39.5%
  • 108353 08/01/03 05:10 S.Woo NMR : 42.28%
  • 108352 08/01/03 04:57 J. Singh controlled access
  • 108347 08/01/03 04:04 S. Woo empty position
  • 108341 08/01/03 03:21 S.Woo multi foils (BeO+12C foils) position
  • 108338 08/01/03 02:53 S.Woo empty position
  • 108337 08/01/03 02:52 S.Woo filled N2 in ref. cell (140 psig) and vented
  • 108331 08/01/03 02:34 J. Singh checked rastered beam spot size
  • 108330 08/01/03 02:30 J. Singh optics with raster ON/OFF
  • 108328 08/01/03 02:29 S. Woo Target moved
  • 108327 08/01/03 02:28 J. Singh MCC called: problem
  • 108320 08/01/03 01:52 S.Woo Multi target(BeO+12C foils) position
  • 108305 08/01/03 00:29 S Woo single target(BeO+12C Long) Position
  • 108296 08/01/03 00:08 averett Swing Shift Summary